WJC:

Chanel deposited $1175 into a savings account that pays a simple annual interest rate of 1.45%. How much interest will she earn after 4 years? • $68.15 • $42.59 • $32.41 • $17.04

jimthompson5910:

You'll use this formula \(\large i = P*r*t\) with i = simple interest earned P = amount deposited (principal) r = interest rate in decimal form t = number of years In this case, i = unknown P = 1175 r = 0.0145 t = 4

WJC:

Is time (4) entered as a exponent?

jimthompson5910:

no, you just multiply out 1175, 0.0145 and 4

jimthompson5910:

1175*0.0145*4 = ??

WJC:

Ok, thank you so very much

jimthompson5910:

No problem

jhonyy9:

good job @jimthompson5910

jimthompson5910:

Thanks I appreciate it